OpenStudy (anonymous):

A cart initially traveling at 10.0 meters per second north accelerates uniformly at 3.0 meters squared to the north for 4.0 seconds. The displacement of the cart from its initial position at the end of this 4.0 seconds is (1) 40 m north (2) 64 m north (3) 88 m north (4) 180 m north I'm taking physics next year but want to be go ahead and learn as much as I can this summer, thank for any help!

OpenStudy (irishboy123):

the equations of motion are your friend: \( \large v = u + at \\ x= ut + \frac{1}{2} a t^2 \\ v^2 = u^2 + 2ax\) use the middle one here.

OpenStudy (irishboy123):

and i think you mean "accelerates uniformly at 3.0 \(m/s^2\)...".